OpenStudy (anonymous):

Drying of Cassava (Tapioca) Root. Tapioca flour is used in many countries for bread and similar products. The flour is made by drying coarse granules of the cassava root containing 66 wt % moisture to 5% moisture and then grinding to produce a flour. How many kg of granules must be dried and how much water removed to produce 5000 kg/h of flour?

OpenStudy (aaronq):

The question tells you that for every 100 grams of raw root, you have 66 grams of water (and consequently 34 grams of pure tapioca flour). If you dry the root until it contains 5% water then you have 34+5=39 grams of flour, for every 100 grams, meaning 39% of the total mass is retained. so we can set up a proportion: \(\sf \dfrac{raw~mass}{100\%}=\dfrac{dried~mass}{39\%}\) we want 5000 kg as the dried mass, so we need: \(\sf \dfrac{raw~mass}{100\%}=\dfrac{5000~kg}{39\%}\rightarrow raw~mass= 12820.5128~kg~\approx 12820.5~ kg \) the difference between the raw mass and the dried mass is the amount of water removed.